If $13,000 is invested at 6% interest compounded monthly, fing the interest earned in 15 years??
Answer by rfer(16322)

A=13000(1+0.06/12)^(12*15)
A=13000(1.005)^180
A=13000(2.454)
A=31,902
---------
31902-13000=18,902.00